OpenStudy (anonymous):

Two lines, R and M, are represented by the following equations: Line R: 2x + 2y = 18 Line M: x + y = 9 Which statement is true about the solution to the set of equations? It is (18, 9). It is (9, 18). There are infinitely many solutions. There is no solution.

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

Okay, let's try and solve the system of equations by using elimination. How many times does the second equation go in the first?

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

Put the equation side by side one another. What do you have to multiply with x to turn into 2x? What do you have to multiply with y to get 2y and what do you have to multiply with 9 to get 18? 2x + 2y = 18 x + y = 9

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Multiply the second equation by 2 then. 2(x + y = 9) It'll equal the first equation. If the second equation and the first equation are equal, what does that mean?

OpenStudy (anonymous):

i dont know

OpenStudy (anonymous):

It means that they're the same line.

OpenStudy (anonymous):

I'm willing to help if you're at least willing to try.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but i don't get what the answer will be

OpenStudy (anonymous):

What do same lines have? Infinite solutions, no solution, or just one solution?

OpenStudy (anonymous):

just 1 solution

OpenStudy (anonymous):

Nope. Try again.

OpenStudy (anonymous):

no solution

OpenStudy (anonymous):

No. They have infinite solutions. Every single point is the same.